The compound known as butylated hydroxytoluene, abbreviated as BHT, contains carbon, hydrogen, and oxygen. A 2.376 g sample of BHT was combusted in an oxygen-rich environment to produce 7.118 g of CO2(g) and 2.331 g of H2O(g).
Insert subscripts to complete the empirical formula of BHT.
1
Expert's answer
2019-10-07T06:42:42-0400

7.118/44 x 12= 1.94127272 g C.

2.331/18 x 2 = 0.259 g H.

2.376-1.94 - 0.259 = 0.177 g O.

1.94/12 : 0.259/1 : 0.177/16 = 0.16166667 : 0.259 : 0.0110625 / 0.011625 = 15 : 24 : 1.

C15H24O.